# Mass Translation and Reflectivity Device Solar Sail Momentum Management Using Predictive Control

### v2

@@ 2.2 Nonlinear System Dynamics
$\sum^{3}_{i=1}\mbf{h}_{b}^{\mathcal{W}_{i}c/a}=\mbf{h}_{b}^{\text{RWs}}+3m_{w}\mbf{r}^{pc^{\times}}_{b}\dot{\mbf{r}}^{pc}_{b}+\sum^{3}_{i=1}\mbf{J}^{\mathcal{W}_{i}c}_{b}{\bm{\omega}}^{ba}_{b},$
@@
$(\underrightarrow{{h}}^{\mathcal{B}c/a})^{{}^{{\raisebox{-2.0pt}{\LARGE$\cdot$}}{\mbox{\scriptsize${a}$}}}}=\underrightarrow{{\tau}}^{\mathcal{B}c},$
@@
$\mbf{J}_{b}^{\mathcal{B}c}\dot{{\bm{\omega}}}_{b}^{ba}+\dot{\mbf{J}}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{J}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}\\ +\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}(\dot{\mbf{r}}_{b}^{ps^{\times}}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{r}_{b}^{ps^{\times}})\dot{\mbf{r}}_{b}^{ps}+\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\mbf{r}_{b}^{ps^{\times}}\ddot{\mbf{r}}_{b}^{ps}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{h}_{b}^{\text{RWs}}+\dot{\mbf{h}}_{b}^{\text{RWs}}={\bm{\tau}}_{b}^{\mathcal{B}c}.$ (5)
Substituting $\dot{\mbf{J}}_{b}^{\mathcal{B}}=-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}(\dot{\mbf{r}}_{b}^{ps^{\times}}\mbf{r}_{b}^{ps^{\times}}+\mbf{r}_{b}^{ps^{\times}}\dot{\mbf{r}}_{b}^{ps^{\times}})$, (5) can be rewritten as
$\mbf{J}_{b}^{\mathcal{B}c}\dot{{\bm{\omega}}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{J}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{h}_{b}^{\text{RWs}}+\dot{\mbf{h}}_{b}^{\text{RWs}}\\ +\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\bigg{(}\mbf{r}_{b}^{ps^{\times}}\ddot{\mbf{r}}_{b}^{ps}+(\dot{\mbf{r}}_{b}^{ps^{\times}}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{r}_{b}^{ps^{\times}})\dot{\mbf{r}}_{b}^{ps}-(\dot{\mbf{r}}_{b}^{ps^{\times}}\mbf{r}_{b}^{ps^{\times}}+\mbf{r}_{b}^{ps^{\times}}\dot{\mbf{r}}_{b}^{ps^{\times}}){\bm{\omega}}_{b}^{ba}\bigg{)}={\bm{\tau}}_{b}^{\mathcal{B}c},$ (6)
@@ 3 MPC-Based Momentum Management Controller
For demonstration purposes in this paper, knowledge of the spacecraft state and the disturbance estimations are assumed to be perfect. Although this assumption is somewhat idealized, it is fair to assume that the solar sail’s onboard flight computer provides an accurate state estimate and that the disturbance torque can be estimated using a Kalman filter, similar to the density estimation approach in [38].
Section 3.1 presents the derivation of the system’s linearized dynamics about the operational state of interest, which is not necessarily an equilibrium point, to facilitate the linear time-invariant (LTI) prediction model for the MPC formulation. The AMT and RCDs exhibit distinct actuation mechanisms that are explicitly accounted for in this framework. The AMT operates continuously, translating the spacecraft bus with a bounded speed, while the RCDs generate discrete on-off torque pulses. Given the large momentum management time step relative to the RWs control time step, improper discretization could lead to significant inaccuracies and misrepresentation of actuator inputs. To better model these momentum management inputs, tailored discretization schemes are employed, with FOH applied to the AMT’s continuous motion and ZOH used for the RCDs’ on-off inputs. The discretization and quantization methodologies are detailed in Sections 3.2 and 3.3. To ensure the RWs maintain their primary role in spacecraft attitude tracking while the momentum management system focuses on mitigating RW angular momentum buildup, a slack variable is introduced in the MPC cost function. This variable penalizes RW angular momentum growth, directing AMT movement and RCD pulses toward momentum management tasks. Additional penalties are applied to minimize AMT motion, as maintaining a fixed AMT position is less costly than a large translation of mass. The complete MPC policy formulation is presented in Section 3.4. While PWM-based quantization is applied to the RCDs’ inputs after the MPC optimization is solved, this introduces a mismatch between the optimized input and the actual applied input to the dynamic propagation. To address this, an iterative backwards-in-time MPC approach is proposed in Section 3.5, which incorporates the effects of quantized RCD torque into the optimization process.
@@ 3.1 Linear Prediction Model
$\delta{\bm{\omega}}^{ba}_{b}=\begin{bmatrix}\delta\omega_{1}\\ \delta\omega_{2}\\ \delta\omega_{3}\end{bmatrix}=\begin{bmatrix}1&0&-\delta\theta_{2}\\ 0&1&\delta\theta_{1}\\ 0&-\delta\theta_{1}&1\end{bmatrix}\begin{bmatrix}\delta\dot{\theta}_{1}\\ \delta\dot{\theta}_{2}\\ \delta\dot{\theta}_{3}\end{bmatrix}$.
@@
${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\big{(}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}+\delta\dot{{\bm{\omega}}}_{b}^{ba}\big{)}+\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big{)}^{\times}\Big{)}\delta\mbf{r}^{ps}_{b}=-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\big{(}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}+{\bar{\mbf{h}}}^{\text{RWs}}_{b}\big{)}\\ +\frac{m_{s}}{m_{p}+m_{s}}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\mbf{f}_{b}^{\text{SRP}}+{\bm{\tau}}^{\text{dist}}_{b}+\delta{\bm{\tau}}^{\text{dist}}_{b}-\dot{{\bar{\mbf{h}}}}_{b}^{\text{RWs}}-\delta\dot{\mbf{h}}_{b}^{\text{RWs}}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\delta\mbf{r}^{ps^{\times}}_{b}+\delta\mbf{r}^{ps^{\times}}_{b}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\big{)}{\bar{\bm{\omega}}}_{b}^{ba}\\ +\frac{m_{s}}{m_{p}+m_{s}}\delta\mbf{r}^{ps^{\times}}_{b}\mbf{f}_{b}^{\text{SRP}}-\delta{\bm{\omega}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\delta{\bm{\omega}}_{b}^{ba}-\delta{\bm{\omega}}_{b}^{ba^{\times}}{\bar{\mbf{h}}}_{b}^{\text{RWs}}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\delta\mbf{h}_{b}^{\text{RWs}}+\delta{\bm{\tau}}^{\text{RCD}}_{b},$ (11)
@@
${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\delta\dot{{\bm{\omega}}}_{b}^{ba}=-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba}\big{)}^{\times}\Big{)}\delta\mbf{r}^{ps}_{b}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big{)}^{\times}\Big{)}\delta\mbf{r}^{ps}_{b}\\ -\frac{m_{s}}{m_{p}+m_{s}}\mbf{f}_{b}^{\text{SRP}^{\times}}\delta\mbf{r}^{ps}_{b}+\Big{(}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+\big{(}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}\big{)}^{\times}+{\bar{\mbf{h}}}_{b}^{\text{RWs}^{\times}}\Big{)}\delta{\bm{\omega}}_{b}^{ba}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\delta\mbf{h}_{b}^{\text{RWs}}\\ -\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big{)}^{\times}\Big{)}\delta\mbf{r}^{ps}_{b}+\delta{\bm{\tau}}^{\text{RCD}}_{b}+\delta{\bm{\tau}}^{\text{dist}}_{b}-\delta\dot{\mbf{h}}_{b}^{\text{RWs}},$ (13)
@@
where $\mbf{x}(t)=\Big{[}\delta{\bm{\theta}}^{\top}\,\,\,\delta{{\bm{\omega}}^{ba}_{b}}^{\top}\,\,\,\delta\mbf{h}^{\text{RWs}^{\top}}_{b}\,\,\,\delta\mbf{e}^{\text{int}^{\top}}\Big{]}^{\top}$, $\mbf{w}(t)=\delta{\bm{\tau}}^{\text{dist}}_{b}$, $\mbf{u}^{\text{AMT}}(t)=\Big{[}\delta r^{\text{AMT}}_{b1}\,\,\,\delta r^{\text{AMT}}_{b2}\Big{]}^{\top}$, $u^{\text{RCD}}(t)=\delta\tau^{\text{RCD}}_{b3}$,
@@
$\frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Big{|}_{\bar{\mbf{x}},\bar{\mbf{u}}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big{(}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba}\big{)}^{\times}\Big{)}-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big{(}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big{)}^{\times}\Big{)}-\frac{m_{s}}{m_{p}+m_{s}}\mbf{f}_{b}^{\text{SRP}^{\times}}\Big{)}$, $\frac{\partial\mbf{f}_{2}}{\partial{\bm{\omega}}_{b}^{ba}}\Big{|}_{\bar{\mbf{x}},\bar{\mbf{u}}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big{(}\big{(}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}\big{)}^{\times}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+{\bar{\mbf{h}}}^{\text{RWs}^{\times}}_{b}-\mbf{K}_{d}\Big{)}$, $\bar{\mbf{x}}=\Big{[}{\bar{\bm{\theta}}}^{\top}\,\,\,{{\bar{\bm{\omega}}}^{ba^{\top}}_{b}}\,\,\,{\bar{\mbf{h}}}^{\text{RWs}^{\top}}_{b}\,\,\,{\bar{\mbf{e}}}^{\text{int}^{\top}}\Big{]}^{\top}$, $\bar{\mbf{u}}=\Big{[}{\bar{\bm{\tau}}}^{\text{dist}^{\top}}_{b}\,\,\,{\bar{\mbf{u}}}^{\text{AMT}^{\top}}\,\,\,\bar{\tau}^{\text{RCD}}_{b3}\Big{]}^{\top}$, and the time-dependencies in the state-space matrices are omitted for brevity. It is worth noting that this is a linear time-varying (LTV) state-space model, where the matrices $\mbf{A}(t)$ and $\mbf{B}(t)=\Big{[}\mbf{B}_{w}(t)\,\,\,\mbf{B}_{u1}(t)\,\,\,\mbf{B}_{u2}(t)\Big{]}$ include time-dependent parameters ${\bar{\bm{\omega}}}_{b}^{ba^{\times}}(t)$, ${\bar{\mbf{h}}}^{\text{RWs}}_{b}(t)$, ${\bar{\mbf{r}}}^{ps}_{b}(t)$, and ${\bar{\bm{\tau}}}^{\text{dist}}_{b}(t)$, which are part of the state ${\bar{\mbf{x}}}(t)$ and input ${\bar{\mbf{u}}}(t)$ to be linearized about. The term ${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}(t)$ is also time-dependent due to the CM translation associated with ${\bar{\mbf{r}}}^{ps}_{b}(t)$.

### v3

@@ 3.2 Nonlinear System Dynamics
$\sum^{3}_{i=1}\mbf{h}_{b}^{\mathcal{W}_{i}c/a}=\mbf{h}_{b}^{\text{RWs}}+3m_{w}\mbf{r}^{pc^{\times}}_{b}\dot{\mbf{r}}^{pc}_{b}+\sum^{3}_{i=1}\mbf{J}^{\mathcal{W}_{i}c}_{b}{\bm{\omega}}^{ba}_{b},$ (8)
@@
$(\underrightarrow{{h}}^{\mathcal{B}c/a})^{{}^{{\raisebox{-2.0pt}{\LARGE$\cdot$}}{\mbox{\scriptsize${a}$}}}}=\underrightarrow{{\tau}}^{\mathcal{B}c},$ (11)
@@
$\mbf{J}_{b}^{\mathcal{B}c}\dot{{\bm{\omega}}}_{b}^{ba}+\dot{\mbf{J}}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{J}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}+\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bm{\omega}}_{b}^{ba^{\times}}\mbf{r}_{b}^{ps^{\times}}\dot{\mbf{r}}_{b}^{ps}\\ +\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\mbf{r}_{b}^{ps^{\times}}\ddot{\mbf{r}}_{b}^{ps}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{h}_{b}^{\text{RWs}}+\dot{\mbf{h}}_{b}^{\text{RWs}}={\bm{\tau}}_{b}^{\mathcal{B}c}.$ (13)
Substituting in $\dot{\mbf{J}}_{b}^{\mathcal{B}}=-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}(\dot{\mbf{r}}_{b}^{ps^{\times}}\mbf{r}_{b}^{ps^{\times}}+\mbf{r}_{b}^{ps^{\times}}\dot{\mbf{r}}_{b}^{ps^{\times}})$ and using the identities $\mbf{u}^{\times}\mbf{v}=-\mbf{v}^{\times}\mbf{u}$ and $\left(\mbf{u}^{\times}\mbf{v}\right)^{\times}=\mbf{u}^{\times}\mbf{v}^{\times}-\mbf{v}^{\times}\mbf{u}^{\times}$ [48], (13) is rewritten as
$\mbf{J}_{b}^{\mathcal{B}c}\dot{{\bm{\omega}}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{J}_{b}^{\mathcal{B}c}{\bm{\omega}}_{b}^{ba}+{\bm{\omega}}_{b}^{ba^{\times}}\mbf{h}_{b}^{\text{RWs}}+\dot{\mbf{h}}_{b}^{\text{RWs}}\\ +\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\bigg(\mbf{r}_{b}^{ps^{\times}}\ddot{\mbf{r}}_{b}^{ps}-2\dot{\mbf{r}}_{b}^{ps^{\times}}\mbf{r}_{b}^{ps^{\times}}{\bm{\omega}}_{b}^{ba}\bigg)={\bm{\tau}}_{b}^{\mathcal{B}c},$ (14)
@@ 4 MPC-Based Momentum Management Controller
For demonstration purposes in this paper, knowledge of the spacecraft state, SRP force and the disturbance estimations are assumed to be perfect. Although this assumption is somewhat idealized, it is fair to assume that the solar sail’s onboard flight computer provides an accurate state estimate and that the disturbance torque can be estimated using a Kalman filter, similar to the density estimation approach in [49].
Section 4.1 presents the derivation of the system’s linearized dynamics about the operational state of interest, which is not necessarily an equilibrium point, to facilitate the linear time-invariant (LTI) prediction model for the MPC formulation. The AMT and RCDs exhibit distinct actuation mechanisms that are explicitly accounted for in this framework. The AMT operates continuously, translating the spacecraft bus with a bounded speed, while the RCDs generate discrete on-off torque pulses. Given the large momentum management time step relative to the RWs control time step, improper discretization could lead to significant inaccuracies and misrepresentation of actuator inputs. To better model these momentum management inputs, tailored discretization schemes are employed, with FOH applied to the AMT’s continuous motion and ZOH used for the RCDs’ on-off inputs. The discretization and quantization methodologies are detailed in Sections 4.2 and 4.3. To ensure the RWs maintain their primary role in spacecraft attitude tracking while the momentum management system focuses on mitigating RW angular momentum buildup, a slack variable is introduced in the MPC cost function. This variable penalizes RW angular momentum growth, directing AMT movement and RCD pulses toward momentum management tasks. Additional penalties are applied to minimize AMT motion, as maintaining a fixed AMT position is less costly than a large translation of mass. The complete MPC policy formulation is presented in Section 4.4. While PWM-based quantization is applied to the RCDs’ inputs after the MPC optimization is solved, this introduces a mismatch between the optimized input and the actual applied input to the dynamic propagation. To address this, an iterative backwards-in-time MPC approach is proposed in Section 4.5, which incorporates the effects of quantized RCD torque into the optimization process.
@@ 4.1 Linear Prediction Model
$\delta{\bm{\omega}}^{ba}_{b}=\begin{bmatrix}\delta\omega_{1}\\ \delta\omega_{2}\\ \delta\omega_{3}\end{bmatrix}=\begin{bmatrix}1&0&-\delta\theta_{2}\\ 0&1&\delta\theta_{1}\\ 0&-\delta\theta_{1}&1\end{bmatrix}\begin{bmatrix}\delta\dot{\theta}_{1}\\ \delta\dot{\theta}_{2}\\ \delta\dot{\theta}_{3}\end{bmatrix}$. (16)
@@
${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\big(\dot{{\bar{\bm{\omega}}}}_{b}^{ba}+\delta\dot{{\bm{\omega}}}_{b}^{ba}\big)+\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big)^{\times}\Big)\delta\mbf{r}^{ps}_{b}\\ =-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\big({\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}+{\bar{\mbf{h}}}^{\text{RWs}}_{b}\big)+\frac{m_{s}}{m_{p}+m_{s}}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\mbf{f}_{b}^{\text{SRP}}+{\bm{\tau}}^{\text{dist}}_{b}+\delta{\bm{\tau}}^{\text{dist}}_{b}-\dot{{\bar{\mbf{h}}}}_{b}^{\text{RWs}}-\delta\dot{\mbf{h}}_{b}^{\text{RWs}}\\ -\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\delta\mbf{r}^{ps^{\times}}_{b}+\delta\mbf{r}^{ps^{\times}}_{b}{\bar{\mbf{r}}}^{ps^{\times}}_{b}\big){\bar{\bm{\omega}}}_{b}^{ba}+\frac{m_{s}}{m_{p}+m_{s}}\delta\mbf{r}^{ps^{\times}}_{b}\mbf{f}_{b}^{\text{SRP}}\\ -\delta{\bm{\omega}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\delta{\bm{\omega}}_{b}^{ba}-\delta{\bm{\omega}}_{b}^{ba^{\times}}{\bar{\mbf{h}}}_{b}^{\text{RWs}}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\delta\mbf{h}_{b}^{\text{RWs}}+\delta{\bm{\tau}}^{\text{RCD}}_{b},$ (20)
@@
${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}\delta\dot{{\bm{\omega}}}_{b}^{ba}=-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba}\big)^{\times}\Big)\delta\mbf{r}^{ps}_{b}\\ -\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big)^{\times}\Big)\delta\mbf{r}^{ps}_{b}\\ -\frac{m_{s}}{m_{p}+m_{s}}\mbf{f}_{b}^{\text{SRP}^{\times}}\delta\mbf{r}^{ps}_{b}+\Big(-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+\big({\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}\big)^{\times}+{\bar{\mbf{h}}}_{b}^{\text{RWs}^{\times}}\Big)\delta{\bm{\omega}}_{b}^{ba}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\delta\mbf{h}_{b}^{\text{RWs}}\\ -\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big)^{\times}\Big)\delta\mbf{r}^{ps}_{b}+\delta{\bm{\tau}}^{\text{RCD}}_{b}+\delta{\bm{\tau}}^{\text{dist}}_{b}-\delta\dot{\mbf{h}}_{b}^{\text{RWs}},$ (22)
@@
where $\mbf{x}(t)=\Big[\delta{\bm{\theta}}^{\top}\,\,\,\delta{{\bm{\omega}}^{ba}_{b}}^{\top}\,\,\,\delta\mbf{h}^{\text{RWs}^{\top}}_{b}\,\,\,\delta\mbf{e}^{\text{int}^{\top}}\Big]^{\top}$, $\mbf{w}(t)=\delta{\bm{\tau}}^{\text{dist}}_{b}$, $u^{\text{RCD}}(t)=\delta\tau^{\text{RCD}}_{b3}$, $\mbf{u}^{\text{AMT}}(t)=\Big[\delta r^{\text{AMT}}_{b1}\,\,\,\delta r^{\text{AMT}}_{b2}\Big]^{\top}$,
@@
$\frac{\partial\mbf{f}_{2}}{\partial\mbf{r}_{b}^{ps}}\Big|_{\bar{\mbf{x}},\bar{\mbf{u}}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(-\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}{\bar{\bm{\omega}}}_{b}^{ba}\big)^{\times}\Big)\\ -\frac{m_{p}^{3}+m_{s}^{3}}{(m_{p}+m_{s})^{2}}\Big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba^{\times}}+\big({\bar{\mbf{r}}}^{ps^{\times}}_{b}\dot{{\bar{\bm{\omega}}}}_{b}^{ba}\big)^{\times}\Big)-\frac{m_{s}}{m_{p}+m_{s}}\mbf{f}_{b}^{\text{SRP}^{\times}}\Big),$
$\frac{\partial\mbf{f}_{2}}{\partial{\bm{\omega}}_{b}^{ba}}\Big|_{\bar{\mbf{x}},\bar{\mbf{u}}}={\bar{\mbf{J}}}_{b}^{\mathcal{B}c^{-1}}\Big(\big({\bar{\mbf{J}}}_{b}^{\mathcal{B}c}{\bar{\bm{\omega}}}_{b}^{ba}\big)^{\times}-{\bar{\bm{\omega}}}_{b}^{ba^{\times}}{\bar{\mbf{J}}}_{b}^{\mathcal{B}c}+{\bar{\mbf{h}}}^{\text{RWs}^{\times}}_{b}-\mbf{K}_{d}\Big),$
$\bar{\mbf{x}}=\Big[{\bar{\bm{\theta}}}^{\top}\,\,\,{{\bar{\bm{\omega}}}^{ba^{\top}}_{b}}\,\,\,{\bar{\mbf{h}}}^{\text{RWs}^{\top}}_{b}\,\,\,{\bar{\mbf{e}}}^{\text{int}^{\top}}\Big]^{\top}$, $\bar{\mbf{u}}=\Big[{\bar{\bm{\tau}}}^{\text{dist}^{\top}}_{b}\,\,\,{\bar{\mbf{u}}}^{\text{AMT}^{\top}}\,\,\,\bar{\tau}^{\text{RCD}}_{b3}\Big]^{\top}$, and the time-dependencies in the state-space matrices are omitted for brevity. It is worth noting that this is a linear time-varying (LTV) state-space model, where the matrices $\mbf{A}(t)$ and $\mbf{B}(t)=\Big[\mbf{B}_{w}(t)\,\,\,\mbf{B}_{u1}(t)\,\,\,\mbf{B}_{u2}(t)\Big]$ include time-dependent parameters ${\bar{\bm{\omega}}}_{b}^{ba^{\times}}(t)$, ${\bar{\mbf{h}}}^{\text{RWs}}_{b}(t)$, ${\bar{\mbf{r}}}^{ps}_{b}(t)$, and ${\bar{\bm{\tau}}}^{\text{dist}}_{b}(t)$, which are part of the state ${\bar{\mbf{x}}}(t)$ and input ${\bar{\mbf{u}}}(t)$ to be linearized about. The term ${\bar{\mbf{J}}}_{b}^{\mathcal{B}c}(t)$ is also time-dependent due to the CM translation associated with ${\bar{\mbf{r}}}^{ps}_{b}(t)$.